Miguez Corporation makes a product with the following standard costs:


Standard Quantity or HoursStandard Price or RateStandard Cost Per Unit
Direct materials
2.3liters$7.00per liter$16.10

Direct labor
0.7hours$22.00per hour$15.40

Variable overhead
0.7hours$2.00per hour$1.40


The company budgeted for production of 2,600 units in September, but actual production was 2,500 units. The company used 5,440 liters of direct material and 1,680 direct labor-hours to produce this output. The company purchased 5,800 liters of the direct material at $7.20 per liter. The actual direct labor rate was $24.10 per hour and the actual variable overhead rate was $1.90 per hour.

The company applies variable overhead on the basis of direct labor-hours. The direct materials purchases variance is computed when the materials are purchased.

1. The variable overhead efficiency variance for September is:

2. The variable overhead rate variance for September is:
Answer #1

Voh efficiency variance

= (SH-AH) * SR

= (2500*0.7 - 1680)*2

= (1750-1680)*2

= 140 Favourable
Answer #3

Correct answer--------------$168 F

Working

Variable Overhead Rate Variance
( Standard Rate - Actual Rate ) x Actual Labor Hours
( $                    2.00 - $                     1.90 ) x 1680
$               168.00
Variance $                 168.00 Favourable-F
Answer #2

Variable overhead rate variance = AH*(AR-SR)

= 1,680 hours*($1.90 per hour – $2 per hour)

= $168 Favorable